OpenStudy (anonymous):

x^2-14x+49=13 will someone check my answer plzzzz

OpenStudy (anonymous):

x=3,-8

OpenStudy (anonymous):

not if it is = 13

OpenStudy (anonymous):

but it all comes out to equalling 13

OpenStudy (anonymous):

perfectly

OpenStudy (anonymous):

Plug 3 into the equation first. 9 -42 + 49 should = 13 ...but it doesn't so no wrong answer

OpenStudy (anonymous):

\[(x-7)^2=13\] \[x-7=\pm\sqrt{13}\] \[x=7\pm\sqrt{13}\]

OpenStudy (anonymous):

so two answers are \[7+\sqrt{13}\] and \[7-\sqrt{13}\]

OpenStudy (anonymous):

Which software are you using SAT

OpenStudy (anonymous):

my brain on this one. lhs is perfect square

OpenStudy (anonymous):

the equation simplifies to x^2-14x+36, now u see that there's no exact integer solution. so use the formula (-b+-sqrt(b^2-4ac))/2a ... this equation gives the quadratic roots for the general equation ax^2+bx+c=0

OpenStudy (anonymous):

but at my disposal i have maple and of course worlfamalpha as it is on line

OpenStudy (anonymous):

Good answer but I like your notations thats what I meant

OpenStudy (anonymous):

but the point of this problem was to recognize the lhs as a perfect square

OpenStudy (anonymous):

Is maple free too?

OpenStudy (anonymous):

oh i am writing latex

OpenStudy (anonymous):

heck no!

OpenStudy (anonymous):

maple cost the earth

OpenStudy (anonymous):

i got it im posting 2 more ? that i need help on if anyone wants to help me

OpenStudy (anonymous):

gianfranco meet me and chat so as not to annoy chloebella. i will show you the latex if you like

OpenStudy (anonymous):

Ok that answers my questions. N ice typing though, well done :)

OpenStudy (anonymous):

Ok